# Automated radiotherapy treatment planning guided by GPT-4Vision

### v2

@@ Abstract
Radiotherapy treatment planning is a time-consuming and potentially subjective process that requires the iterative adjustment of model parameters to balance multiple conflicting objectives. Recent advancements in frontier AI models offer promising avenues for addressing the challenges in planning and clinical decision-making. This study introduces GPT-RadPlan, an automated treatment planning framework that integrates radiation oncology knowledge with the reasoning capabilities of large multi-modal models, such as GPT-4Vision (GPT-4V) from OpenAI. Via in-context learning, we incorporate clinical protocols for various disease sites to enable GPT-4V to acquire treatment planning domain knowledge. The resulting GPT-RadPlan agent is integrated into our in-house inverse treatment planning system through an API. For a given patient, GPT-RadPlan acts as both plan evaluator and planner, first assessing dose distributions and dose-volume histograms (DVHs), and then providing “textual feedback” on how to improve the plan. In this manner, the agent iteratively refines the plan by adjusting planning parameters, such as weights and objective doses, based on its suggestions. The efficacy of the automated planning system is showcased across multiple prostate and head & neck cancer cases, where we compared GPT-RadPlan results to clinical plans produced by human experts. In all cases, GPT-RadPlan either outperformed or matched the clinical plans, demonstrating superior target coverage and organ-at-risk sparing. Consistently satisfying the dosimetric objectives in the clinical protocol, GPT-RadPlan represents the first multimodal large language model agent that mimics the behaviors of human planners in radiation oncology clinics, achieving promising results in automating the treatment planning process without the need for additional training.
@@ 1 Introduction
Treatment planning of modern radiation therapy (RT) modalities, such as intensity modulated RT (IMRT) and volumetric modulated arc therapy (VMAT) is an inverse process solving an optimization problem based on dose prescriptions Oelfke and Bortfeld (2001); Xing and Chen (1996); Yu and Tang (2011). This optimization problem is typically defined as a constrained optimization minimizing a weighted cost function that balances multiple conflicting objectives Webb (2003); Xing et al. (1999); Yang and Xing (2004). The result is greatly impacted by the value of the optimization parameters (e.g., the objective weights in the cost function), often resulting in plans that cannot meet all objectives at once.
@@
A main challenge in treatment planning is translating the overall clinical goals into weighted objective functions and dose constraints yielding acceptable plans. Human planners resort to a trial-and-error approach, whereby they iteratively adjust the optimization parameters based on the results from the optimization problem until the resulting plans meet the clinical requirements Xing et al. (1999). Treatment planning is potentially subjective, since changes in the treatment parameters are based on the experience of the planner Hussein et al. (2018). Additionally, trial-and-error involves repeatedly using the computationally-expensive optimization algorithms over many iterations, with potentially multiple interactions between planner and radiation oncologists in between, making the entire process time-consuming and costly, and presenting a bottleneck in precision RT.
@@
Multiple approaches such as knowledge-based planning or multi-criteria optimization have been investigated to reduce the load on human planners Hussein et al. (2018). Knowledge-based approaches utilize prior experience to improve plans or derive good starting points for subsequent trial-and-error adjustment. Some examples include selecting closest matching patients Petrovic et al. (2016); McIntosh and Purdie (2016); Yang and Xing (2004); Huang et al. (2023), or using models that predict achievable dose volume histograms (DVHs) Munter and Sjölund (2015). Alternatively, the multi-criteria approach often involves a “wish list” containing objective functions with assigned priorities, which iteratively steers planning to meet as many objectives as possible ordered by their priority Breedveld et al. (2012, 2007); Voet et al. (2013, 2012).
@@
With the recent success of deep learning, multiple neural network-based approaches have also been proposed to assist treatment planning, mostly based on U-net Ronneberger et al. (2015) or generative adversarial frameworks Goodfellow et al. (2014). Most works focus on predicting dose distributions that may be achievable based on previous clinical plans. As an input, these models take organ masks Ma, K. Buyyounouski, et al. (2019); Fan et al. (2019); Nguyen et al. (2019); Ma, Kovalchuk, et al. (2019), possibly also including beam information Barragán-Montero et al. (2019). Furthermore, these learning-based methods require large datasets and extensive training to ensure generalization across disease sites and device settings, making it difficult to integrate them into clinical workflow.
@@
Adjusting the optimization parameters can be defined as a decision-making problem that can ideally be fully automated. Framing treatment planning as a sequence of actions taken by a planner, reinforcement learning (RL) has been recently applied to mimic human planners in RT workflows Xu et al. (2022). RL methods learn in a trial-and-error manner how to take actions maximizing a cumulative reward (i.e., a scalar based on dose metrics or target conformity). Some initial studies demonstrate that RL agents can learn to modify optimization parameters such as weights and objective doses Shen et al. (2020), while other works train agents that directly optimize machine parameters Hrinivich and Lee (2020). In practice, RL-based approaches face multiple problems. First, designing reward functions of RL that balance between different objectives — also known as the “reward design problem” - is notoriously difficult because agents are susceptible to reward hacking Norvig and Intelligence (2002); Wiering and Van Otterlo (2012). Indeed, it is not straightforward how to write a reward function to evaluate the treatment plan that requires balancing multiple clinical objectives. While RL agents can learn from labeled examples, this is not possible with a single example; RL agents need to gather large amounts of labeled data to capture the nuances of different users’ preferences and objectives, which has shown to be costly Zhang et al. (2016). Additionally, both approaches do not generalize well to new users or new environments that have different objectives, necessitating re-design or re-collection of labeled data.
@@
Our aim is to leverage multi-modal LLMs to create an easier way for RT professionals to communicate their preferences and to facilitate the inverse treatment planning process. In the past few years, LLMs that are trained on internet-scale data have witnessed significant advancements and shown an impressive ability to learn in-context from a few examples S. Liu et al. (2023). Our key insight is that the scale of data that multi-modal LLMs have been trained on makes them great in-context learners S. Liu et al. (2023), also allowing them to capture meaningful commonsense priors about human behavior as well as expertise on specific domains. Given a few examples and/or a description demonstrating the radiation oncologist’s objective, an LLM should be able to provide an accurate evaluation of a new test example and provide solutions for further improvement.
@@
In this work, we explore how to prompt multi-modal LLMs as a proxy reward function (critic) and agent (actor) for radiotherapy treatment planning. Specifically, we propose a novel framework to fully automate planning using pre-trained multi-modal LLMs. The proposed GPT-RadPlan is capable of learning to mimic human planners without any domain-specific training, using only three clinical plans as a reference, in the form of text, DVH tables, and dose distribution images. GPT-RadPlan iteratively adjusts optimization parameters, directly assessing whether the plan meets the clinical requirements based on agreement with the clinical protocol. In this case, GPT-RadPlan acts both as an evaluator and planner, first determining which aspects of the plan can be improved on the clinical requirements, and subsequently suggesting new optimization weights to address the identified deficiencies. Through our experiments, we compare GPT-RadPlan’s results to clinically approved prostate cancer and head & neck plans. There are two advantages to prompting multi-modal LLMs for treatment planning: (1) we can leverage LLM’s in-context learning abilities and prior knowledge of treatment planning so that users only need to provide a handful of examples of desirable plans and (2) we can specify our preferences intuitively using language.
@@
Figure 1: Overview of GPT-RadPlan workflow. a. Integration of multi-modal LLM-based RT treatment planning into the existing clinical workflow, using GPT-4V. GPT-RadPlan plan can communicate with the users in human language and iteratively adjusts optimization weights. b. Modules for the application of GPT-4V at different steps of the RT workflow. Based on the independent evaluation of the dose distribution by an image AI expert, and the DVH assessment from the DVH AI expert, GPT-4V reviews and/or approves the plan, providing feedback to the AI planner. The planner produces a new set of parameters based on the historical trajectory and feedback on the current plan.
@@ 2.1 Problem formulation
We opt for a two-loop optimization formulation Xing et al. (1999); Huang et al. (2022) consisting of (i) an inner loop inverse planning step and (ii) an outer loop step optimizing the parameters of the inner loop. As such, the inner loop performs traditional fluence map optimization, trying to recover the optimal fluence map $x$ by minimizing a cost function composed of multiple weighted objectives for different targets and organs at risk, defined as
$\min_{x}$ $\sum_{t}w_{t}([Kx]_{t}-d_{t})^{2}+\sum_{s}w_{s}\Theta([Kx]_{s}-d_{s})([Kx]_{s}-d_{s}),$ (1)
s.t. $D_{95}([Kx]_{t})=d_{t}$
@@
Here, $\{w_{t}\}_{t=1}^{N_{t}}$ and $\{w_{s}\}_{s=1}^{N_{s}}$ are the weights that balance the different objectives for $N_{t}$ and $N_{s}$ different PTV targets and OARs, respectively. $K$ represents the dose influence matrix, containing the dose per unit fluence delivered to voxels in the volume by beamlets, and $\{d_{t}\}_{t=1}^{N_{t}}$ and $\{d_{s}\}_{s=1}^{N_{s}}$ are the scalar objective doses for each structure. Essentially, the whole cost function penalizes squared deviations from the target objective doses, regardless of the positive or negative sign, while the OAR terms only penalize squared over-dosing for doses above $d_{s}$. As such, $\Theta$ represents the Heaviside function constraining the objective to only positive values. The minimization is constrained for only positive $x$ values, and forcing that the $D_{95}$ — the maximum common dose received by 95% of the structure volume — equals the prescribed dose for the targets.
We design GPT-RadPlan to operate the outer loop, finding the optimal target and OAR weights that are used to minimize Equation 1. GPT-RadPlan consists of three algorithmic components: a) environment as the context that leverages GPT-4V prior knowledge, b) vision expert that analyzes the dose distribution images, and c) DVH expert that analyzes the DVH tables and dose statistics.
@@ 2.2 Overall framework
As shown in Figure 1, GPT-RadPlan introduces LLM-based agents as the evaluator and planner. The system incorporates three fundamental modules to facilitate the treatment planning process, enabling iterative reasoning, planning, and continuous interaction between the agents and the treatment planning system. The task concludes either when the clinical protocols are successfully achieved or when the maximum iteration limit is reached. The functionalities of each module are as follows:
@@
The evaluator assumes a central role within the workflow of GPT-RadPlan. It receives the present dose distribution images, DVHs, and dose statistics, enabling evaluation and learning from the planner’s historical trajectories under the context of clinical protocols. Functioning as a centralized coordinator, the evaluator engages in reasoning and providing suggestions on directions of improvement on the current plant. Such suggestions from the critic are then used to adjust the optimization parameters and then optimize the plan, allowing for a continuous cycle of plan improvement.
To capture dose distribution images and DVHs of the treatment plan, we employ GPT-4V as a vision expert on dose distribution images and DVHs. To assist the vision expert in interpreting these images, we request GPT-4V to analyze the regional color intensity and distribution around the PTV and OARs. On the other hand, the DVH expert will take the DVH and related dose statistics as input, the DVH is first transformed into a table and then translated into a text sequence. The corresponding clinical protocol with prescribed dose and clinical requirements is also provided, so the GPT-4V can perform reasoning on the DVH based on the protocol. Aggregating the outputs of the vision expert and DVH expert, the GPT-4V provides final feedback on how to improve the current treatment plan.
For a treatment plan, we obtain its dose distribution images $I$ and the DVH $\delta$, the evaluator then evaluates them to derive feedback $f$ based on the prompt $p$ as:
$f=g_{\text{evaluator}}(p;I,\delta).$
@@
The memory module serves to store crucial information needed during the planning process to aid the accumulation of useful knowledge and enhance the agent’s decision-making capabilities. Specifically, the memory module is used to store the historical trajectory of the plans and the corresponding optimization parameters. The memory module also incorporates a mechanism for filtering redundant information. If the planning process is long-term, it retains only the most recent $L$ steps of dose distribution images, DVHs, feedback and the optimization parameters $<(I_{t-L+1},\delta_{t-L+1},f_{t-L+1},a_{t-L+1}),\dots,(I_{t},\delta_{t},f_{t},a_{t})>$. This assists the agent in comprehending the relationship between the optimization parameters and changes in the treatment plan.
@@
The planning module fulfills the vital function of converting the text-based feedback from the evaluation module to optimization parameters that can be processed by the treatment planning system. Such parameters can be beam angles for IMRT or target and OAR weights. The planning module receives the present treatment plan, and feedback from the evaluation module and extracts pertinent details from the memory module, enabling learning from its historical trajectories. Furthermore, a few examples of paired treatment plans and optimization parameters are provided to further facilitate the planning module’s understanding of the relationship between the optimization parameters and changes in the treatment plan $a=g_{\text{planning}}(p;I_{t},\delta_{t},f_{t})$.

### v3

@@ Abstract
Objective: Radiotherapy treatment planning is a time-consuming and potentially subjective process that requires the iterative adjustment of model parameters to balance multiple conflicting objectives. Recent advancements in frontier Artificial Intelligence (AI) models offer promising avenues for addressing the challenges in planning and clinical decision-making. This study introduces GPT-RadPlan, an automated treatment planning framework that integrates radiation oncology knowledge with the reasoning capabilities of large multi-modal models, such as GPT-4Vision (GPT-4V) from OpenAI.
Approach: Via in-context learning, we incorporate clinical requirements and a few (3 in our experiments) approved clinical plans with their optimization settings, enabling GPT-4V to acquire treatment planning domain knowledge. The resulting GPT-RadPlan system is integrated into our in-house inverse treatment planning system through an application programming interface (API). For a given patient, GPT-RadPlan acts as both plan evaluator and planner, first assessing dose distributions and dose-volume histograms (DVHs), and then providing “textual feedback” on how to improve the plan to match the physician’s requirements. In this manner, GPT-RadPlan iteratively refines the plan by adjusting planning parameters, such as weights and dose objectives, based on its suggestions.
Main results: The efficacy of the automated planning system is showcased across 17 prostate cancer and 13 head & neck cancer VMAT plans with prescribed doses of 70.2 Gy and 72 Gy, respectively, where we compared GPT-RadPlan results to clinical plans produced by human experts. In all cases, GPT-RadPlan either outperformed or matched the clinical plans, demonstrating superior target coverage and reducing organ-at-risk doses by 5 Gy on average (15% for prostate and 10-15% for head & neck).
Significance: Consistently satisfying the the dose-volume objectives in the clinical protocol, GPT-RadPlan represents the first multimodal large language model agent that mimics the behaviors of human planners in radiation oncology clinics, achieving promising results in automating the treatment planning process without the need for additional training.
@@ 1 Introduction
Treatment planning of modern radiation therapy (RT) modalities, such as intensity modulated RT (IMRT) and volumetric modulated arc therapy (VMAT) is an inverse process solving an optimization problem based on dose prescriptions [1, 2, 3]. This optimization problem is typically defined as a constrained optimization minimizing a weighted cost function that balances multiple conflicting objectives [4, 5, 6]. The result is greatly impacted by the value of the optimization parameters (e.g., the objective weights in the cost function), often resulting in plans that cannot meet all objectives at once.
@@
A main challenge in treatment planning is translating the overall clinical goals into weighted objective functions and dose constraints yielding acceptable plans. Human planners resort to a trial-and-error approach, whereby they iteratively adjust the optimization parameters based on the results from the optimization problem until the resulting plans meet the clinical requirements [5]. Treatment planning is potentially subjective, since changes in the treatment parameters are based on the experience of the planner [7]. Additionally, trial-and-error involves repeatedly using the computationally-expensive optimization algorithms over many iterations, with potentially multiple interactions between planner and radiation oncologists in between, making the entire process time-consuming and costly, and presenting a bottleneck in precision RT.
@@
Multiple approaches such as knowledge-based planning or multi-criteria optimization have been investigated to reduce the load on human planners [7]. Knowledge-based approaches utilize prior experience to improve plans or derive good starting points for subsequent trial-and-error adjustment. Some examples include selecting closest matching patients [8, 9, 6, 10], or using models that predict achievable dose volume histograms (DVHs) [11]. Alternatively, the multi-criteria approach often involves a “wish list” containing objective functions with assigned priorities, which iteratively steers planning to meet as many objectives as possible ordered by their priority [12, 13, 14, 15].
@@
With the recent success of deep learning, multiple neural network-based approaches have also been proposed to assist treatment planning, mostly based on U-net [16] or generative adversarial frameworks [17]. Most works focus on predicting dose distributions that may be achievable based on previous clinical plans. As an input, these models take organ masks [18, 19, 20, 21], possibly also including beam information [22]. Furthermore, these learning-based methods require large datasets and extensive training to ensure generalization across disease sites and device settings, making it difficult to integrate them into clinical workflow.
@@
Adjusting the optimization parameters can be defined as a decision-making problem that can ideally be fully automated. Framing treatment planning as a sequence of actions taken by a planner, reinforcement learning (RL) has been recently applied to mimic human planners in RT workflows [23]. RL methods learn in a trial-and-error manner how to take actions maximizing a cumulative reward (i.e., a scalar based on dose metrics or target conformity). Some initial studies demonstrate that RL agents can learn to modify optimization parameters such as weights and objective doses [24], while other works train agents that directly optimize machine parameters [25]. In practice, RL-based approaches face multiple problems. First, designing reward functions of RL that balance between different objectives — also known as the “reward design problem” - is notoriously difficult because agents are susceptible to reward hacking, which is when they exploit flaws or loopholes in the reward function to achieve high rewards in unintended ways, rather than learning the desired behavior. [26, 27]. Indeed, it is not straightforward how to write a reward function to evaluate the treatment plan that requires balancing multiple clinical objectives. While RL agents can learn from labeled examples, this is not possible with a single example; RL agents need to gather large amounts of labeled data to capture the nuances of different users’ preferences and objectives, which has shown to be costly [28]. Additionally, both learning and RL based methods do not generalize well to new users or new environments that have different objectives, necessitating re-design or re-collection of labeled data.
@@
Our aim is to leverage multi-modal large language models (LLMs) to create an easier way for RT professionals to communicate their preferences and to facilitate the inverse treatment planning process. In the past few years, LLMs that are trained on internet-scale data have witnessed significant advancements and shown an impressive ability to learn in-context from a few examples [29]. Our key insight is that the scale of data that multi-modal LLMs have been trained on makes them great in-context learners [29], also allowing them to capture meaningful commonsense priors about human behavior as well as expertise on specific domains. Given a few examples and/or a description demonstrating the radiation oncologist’s objective, an LLM should be able to provide an accurate evaluation of a new plan and provide solutions for further improvement.
@@
In this work, we explore how to prompt multi-modal LLMs as a proxy reward function (evaluator) and agent (planner) for radiotherapy treatment planning. Specifically, we propose a novel framework to fully automate the planning process using pre-trained multi-modal LLMs. The proposed GPT-RadPlan is capable of learning to mimic human planners without any domain-specific training, using only three clinical plans as a reference, in the form of optimized weights, DVH tables (the DVH in a tabular format where each row contains doses received for a given fraction of the volume of each of the organs), and dose distribution images. GPT-RadPlan iteratively adjusts optimization parameters, directly assessing whether the plan meets the clinical requirements based on agreement with the clinical protocol. In this case, GPT-RadPlan acts both as an evaluator and planner, first determining which aspects of the plan can be improved on the clinical requirements, and subsequently suggesting new optimization weights to address the identified deficiencies. Through our experiments, we compare GPT-RadPlan’s results to clinically approved treatment plans for prostate and head & neck cancer (HN) patients. There are two advantages to prompting multi-modal LLMs for treatment planning: (1) we can leverage LLM’s in-context learning abilities and prior knowledge of treatment planning so that users only need to provide a handful of examples of desirable plans, and (2) we can specify our preferences intuitively using human language in text.
@@
Figure 1: Overview of GPT-RadPlan workflow. a. Integration of multi-modal LLM-based RT treatment planning into the existing clinical workflow, using GPT-4V. GPT-RadPlan can be human-in-the-loop where the users can use human language to provide feedback on how to improve the current plan when needed, and iteratively adjust optimization weights. b. Modules for the application of GPT-4V at different steps of the RT workflow. Based on the independent evaluation of the dose distribution by an image AI expert, and the DVH assessment from the DVH AI expert, GPT-4V reviews the plan, providing feedback to the AI planner, and approves the plan after the plan is good enough. The planning module translates feedback from the evaluation module into concrete optimization parameters, enabling iterative refinement of the treatment plan. GPT-4V’s in-context learning capabilities facilitate this process without requiring additional model training, relying instead on input prompts and historical planning data.
@@ 2.1 Problem formulation
We opt for a two-loop optimization formulation [5, 30] consisting of (i) an inner loop performing inverse fluence optimization and (ii) an outer loop step optimizing the parameters of the inner loop. As such, the inner loop performs traditional fluence map optimization, trying to obtain the optimal fluence map $x$ by minimizing a cost function composed of multiple weighted objectives for different targets and organs at risk, defined as
$\min_{x}$ $\sum_{p}w_{p}([Ax]_{p}-d_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}})^{2}+\sum_{s}w_{s}\Theta([Ax]_{s}-d_{s})([Ax]_{s}-d_{s}),$ (1)
s.t. $D_{95}([Ax]_{p})=d_{p}$
@@
Here, $\{w_{p}\}_{p=1}^{N_{p}}$ and $\{w_{s}\}_{s=1}^{N_{s}}$ are the weights that balance the different objectives for $N_{p}$ PTV targets and $N_{s}$ organs at risk (OARs), respectively. $A$ represents the dose influence matrix, containing the dose per unit fluence delivered to voxels in the volume by beamlets, and $\{d_{p}\}_{p=1}^{N_{p}}$ and $\{d_{s}\}_{s=1}^{N_{s}}$ are the scalar objective doses for each structure. Essentially, the whole cost function penalizes squared deviations from the target objective doses, regardless of the positive or negative sign, while the OAR terms only penalize squared over-dosing for doses above $d_{s}$. As such, $\Theta$ represents the Heaviside function constraining the objective to only positive values. The minimization is constrained for only positive $x$ values, and forcing that the $D_{95}$the minimum dose that covers at least 95% og the target volume — equals the prescribed dose for the targets.
We design GPT-RadPlan to operate the outer loop, finding the optimal target and OAR weights that are used to minimize Equation 1. GPT-RadPlan consists of three algorithmic components: a) an evaluation module that leverages GPT-4V prior knowledge as well as reference plans to evaluate the plan in the current iteration, b) a memory module that stores the plans and the corresponding optimization parameters in the previous iterations, c) a planning module that converts the text-based feedback from the evaluation module to PTV and OAR weights for inverse planning optimizer.
@@ 2.2 Overall framework
As shown in Figure 1, GPT-RadPlan introduces LLM-based agents as the evaluator and planner. The system incorporates three fundamental modules to facilitate the treatment planning process, enabling iterative reasoning, planning, and continuous interaction between the agents and the treatment planning system. The task concludes either when the patient-specific requirements in the physician’s intent are successfully achieved or when the maximum iteration limit is reached. Algorithm 1 illustrates the detailed algorithm which breakdowns the GPT-RadPlan step-by-step. The functionalities of each module are as follows:
Algorithm 1 GPT-RadPlan Auto-Plan Framework
Input: Patient data ($P$), Reference plans ($R$), Clinical requirements ($C$), Max iterations ($N$)
Output: Optimized parameters $w$
1:Initialize EvaluationModule with reference cases and clinical objectives
2:Initialize MemoryModule for storing historical trajectories
3:Initialize PlanningModule for generating optimization parameters
4:Extract initial dose distribution $I_{0}$ and DVH table $\delta_{0}$ from patient data
5:for $i=1$ to $N$ do
6:    Feedback $F_{i}\leftarrow\textbf{EvaluationModule}.\text{evaluate}(I_{i-1},\delta_{i-1})$triangleright Evaluate current plan
7:    MemoryModule.store($I_{i-1},\delta_{i-1},F_{i}$) $\triangleright$ Store the current state in memory
[…]
9:    $I_{i},\delta_{i}\leftarrow\text{innerLoopOptimization}(w)$triangleright Generate new plan using updated parameters
[…]
11:         break $\triangleright$ Stop if the plan meets clinical objectives
12:    end if
13:end for
14:return $w$
@@
The evaluator assumes a central role within the workflow of GPT-RadPlan, assessing the level of agreement between the current plan and the physician’s intent. It receives the present dose distribution images, DVHs, and dose statistics such as the precomputed minimum, maximum, and mean doses based on DVHs together with the physician’s intent (in textual form as a list of requirements). Functioning as a centralized coordinator, the evaluator engages in reasoning and providing suggestions on directions of improvement for the current plant. Such suggestions from the evaluation module are then used to adjust the optimization parameters and then optimize the plan, allowing for a continuous cycle of plan improvement.
To properly evaluate both qualitative aspects about the dose distribution and quantitative values and DVHs of the treatment plan, we separately employ GPT-4V as a vision and DVH expert. The vision expert interprets the dose distribution images, analyzing the regional color intensity based on the color bar, presence of hot or cold spots, and distribution and spillage around the PTV and OARs. This analysis is done using dose images were automatically cropped around each structure, using the contours with a 5-pixel margin, as well as the full dose distribution in the sagittal, coronal, and axial axes.
On the other hand, the DVH expert will take the DVH and related dose statistics as input. The DVH is first transformed into a table and then translated into a text sequence. The corresponding physician’s intent with prescribed dose and clinical requirements for the patient is also provided. The model then reasons about the level of agreement of the DVH and metrics with the intent. Aggregating the outputs of the vision expert and DVH expert, the GPT-4V provides final feedback on how to improve the current treatment plan.
For a treatment plan, we obtain its dose distribution images $I$ and the DVH $\delta$, the evaluator then evaluates them to derive feedback $F$ based on the prompt as $F=\text{EvaluationModule}(I,\delta).$
@@
The memory module serves to store crucial information needed during the planning process to aid the accumulation of useful knowledge and enhance the agent’s decision-making capabilities. Specifically, the memory module is used to store the plans and the corresponding optimization parameters from past iterations for the same patient. The memory module also incorporates a mechanism for filtering redundant information. If the planning process takes too many iterations, it retains only the most recent $L$ steps of dose distribution images, DVHs, feedback and the optimization parameters, forming a set in the order of steps $\{(I_{t-L+1},\delta_{t-L+1},F_{t-L+1},w_{t-L+1}),\dots,(I_{t},\delta_{t},F_{t},w_{t})\}$. This assists the agent in comprehending the relationship between the optimization parameters and changes in the treatment plan.
@@
The planning module fulfills the vital function of converting the text-based feedback from the evaluation module to optimization parameters that can be processed by the treatment planning system. The planning module receives the current iteration of the treatment plan and the feedback from the evaluation module and extracts pertinent details from the memory module, enabling learning from its historical trajectories.
To help with initialization and guide the parameter adjustment process, the planning module receives some reference examples of previously approved treatment plans from the same treatment site and clinical protocol, together with the optimization parameters used to obtain these plans. This further facilitates the planning module’s understanding of the relationship between the optimization parameters and changes in the treatment plan. As a result, the model outputs the optimization parameters that will drive the inner optimization loop of the next iteration $w=\text{PlanningModule}(I_{t},\delta_{t},F_{t})$. The example prompts can be found in Section A of the supplementary.